2. A company made the following merchandise purchases and sales during the current month:

July 1 purchased 500 units at $15 each

July 5 purchased 300 units at $20 each

July 14 purchased 300 units at $25 each

July 20 purchased 250 units at $30 each

July 30 sold 575 units at $80 each

There was no beginning inventory. What is the Cost of Goods Sold under FIFO? What is the Cost of Goods Sold under LIFO? What is the Cost of Goods Sold under Weighted Average?
Answer #1

Answer :

Calculation of cost of goods sold

Using FIFO

Unit Sold = 575

Cost of goods sold

July 1 cost of 500 unit = 500*15 =$7500

July 14 Cost of 75 Unit = 75*25 =$1875

Total Cost = $9375 Answer

Using LIFO

Unit Sold =575

Cost of goods sold  

July 20 Cost of 250 Unit =250*30 =$7500

July 14 cost of 300 unit =300*25 =$7500

July 5 Cost of 25 (575-550)unit =$500 (25*20)

Total =$15,500/- Answer

Using Weighted Average Method

Unit Sold = 575

Cost of goods sold = Unit sold * Weighted average cost of per unit

Cost of goods sold = 575*21.11

Cost of goods sold = $21,138.25/- Answer

Calculation of Weighted Average cost

= Total cost / Total no of unit

=$28500/1350

=21.11/- per unit

Date of purchase Unit Rate Total
Jily 1 500 $15 $7500
July 5 300 $20 $6000
July 14 300 $25 $7500
July 20 250 #30 $7500
Total 1350 $28500